OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4).

OpenStudy (anonymous):

g(x) = x2 + 3 g(4) = 4`2 +3 g(4) = (4*4) + 3 g(4) = 16 + 3 =19 you just substitute x = 4

OpenStudy (anonymous):

options are: 11 19 16 8

OpenStudy (♪chibiterasu):

19 As ayo said.

OpenStudy (anonymous):

no Elsa.it's x to the power of 2